Which of the following would have the greatest dipole? Why?

a) Cl—Br

b) N—F

c) O—F

d) P—S
Answer #1

Electronegativity of Cl = 3.16

Electronegativity of Br = 2.96

Electronegativity difference Cl-Br = 3.16 - 2.96 = 0.2

Electronegativity of N = 3.04

Electronegativity of F = 3.98

Electronegativity difference N-F = 3.98 - 3.04  = 0.94

Electronegativity of O = 3.44

Electronegativity of F = 3.98

Electronegativity difference O-F = 3.98 - 3.44 = 0.54

Electronegativity of P = 2.19

Electronegativity of S = 2.58

Electronegativity difference in P-S = 2.58 - 2.19 = 0.39

Among all, the electrinegativity difference in N-F bond is higher. If electronegaticvity difference is higher the it sould have greatest dipole.

Hence N-F would have the greatest dipole.

Option B is correct answer